OpenStudy (anonymous):

Rationalize the denominator. -9/2sqareroot10 I have the answer, but I need help with the process.

OpenStudy (anonymous):

\[\frac{-9}{2\sqrt{10}}=\frac{-9}{2\sqrt{10}} \times \frac{\sqrt{10}}{\sqrt{10}}=\frac{-9\sqrt{10}}{2\times10}=\frac{-9\sqrt{10}}{20}\]

OpenStudy (anonymous):

Thank you!